OpenStudy (anonymous):

True or False: There are values of t so that the equation cos t =5/3 holds There are values of t so that sin t = 0.3 and cos t = - 0.6

ganeshie8 (ganeshie8):

hint : the value of cos(t) lies between -1 and 1

ganeshie8 (ganeshie8):

\(\large -1 \le \cos(t) \le 1\)

OpenStudy (anonymous):

\[\sin ^2t+\cos ^2t=1\]

OpenStudy (anonymous):

So, are they both false?

OpenStudy (anonymous):

correct